# Symplectic Lattices and GKP Codes - Simple Randomized Constructions from Cryptographic Lattices

### v1

@@ 8.2 Symplectic lattices from R-SIS lattices
$\bar{I}_{n-1}:=\left(\begin{matrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\vdots&0&0\end{matrix}\right),\quad\sigma_{n}:=\left(\begin{matrix}1&0\\ 0_{n}&-\bar{I}_{n-1}\end{matrix}\right).$
@@
$M_{\textbf{h}}:=M_{\rho(\textbf{h})}\quad\text{and}\quad\bar{M}_{\textbf{h}}:=M_{\bar{\rho}(\textbf{h})}.$
@@ 8.3 Symplectic lattices from M-SIS lattices
that is, we replace entries $\textbf{h}_{ij}$ by the matrix representations of multiplication with $\textbf{h}_{ij}.$ Consider a symmetric matrix H over $R$. Matrix $M_{\rho(\textbf{H})}$ is not symmetric, since matrix representations of multiplications with ring elements are not symmetric. However, for symmetric H the matrix
@@
$M_{\textbf{H}}:=M_{\rho(\textbf{H})}\quad\text{and}\quad\bar{M}_{\textbf{H}}:=M_{\bar{\rho}(\textbf{H})}.$
@@ 9.1 Analyzing the probability
$X^{n}+1:=\prod_{j=1}^{\ell}p_{j}(X)$ (17)
be the factorization of $X^{n}+1$ into irreducible factors modulo $q$ (see (11). Later, using (12) and Theorem 21, we consider this factorization in more detail. For now, set ${\cal I}:=\{1,\ldots,\ell\}$ and for $\textbf{z}\in R^{2k}$
$\bar{{\cal I}}_{\textbf{z}}:=\{j\in{\cal I}:\forall h\textbf{z}_{1h}=\textbf{z}_{2h}=0\mod p_{j}(X)\}\quad\text{and}\quad{\cal I}_{\textbf{z}}:={\cal I}\setminus\bar{{\cal I}}_{\textbf{z}}.$ (18)
Set
$d_{\textbf{z}}:=\sum_{j\in{\cal I}_{\textbf{z}}}\deg(p_{j}(X)),\bar{d}_{\textbf{z}}:=n-d_{\textbf{z}}.$
@@
$\textbf{z}\in{\cal L}^{\perp}(\textbf{H})\Leftrightarrow\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod qR\Leftrightarrow\forall j\in{\cal I}\;\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{j}(X).$
[…]
$\Pr[\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{j}(X);\textbf{H}\leftarrow\mathbb{F}_{q^{d_{j}}}^{k\times k}\;\text{symmetric}]=\frac{1}{q^{d_{j}}}.$
For distinct elements in ${\cal I}_{\textbf{z}}$ these events are independent. Observing
$d_{\textbf{z}}=\sum_{j\in{\cal I}_{\textbf{z}}}d_{j}$
proves the lemma. ∎
@@ 9.2 Bounds on the number of short vectors
Let $\cal T$$\subset{\cal I}$. In this subsection we derive bounds on the number of short elements in $\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2})\in R^{2k}$ satisfying
$\forall j\not\in{\cal T}:\textbf{z}_{1}=\textbf{z}_{2}=0\mod p_{j}(X).$
To state this more precisely, for ${\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset$, set
  • $p_{\cal T}(X):=\prod_{l\in T}p_{j}(X),\quad\bar{p}_{\cal T}(X):=\prod_{l\not\in T}p_{j}(X)$
  • $d_{\cal T}:=\deg(p_{\cal T}(X)),\quad\bar{d}_{\cal T}:=\deg(\bar{p}_{\cal T}(X))$
@@
  • $Z_{\cal T}:=\left\{\textbf{z}\in R^{2k}:\textbf{z}_{1}=\textbf{z}_{2}=0\mod\bar{p}_{\cal Z}(X)\right\}.$
@@
Properties 1. and 2. are immediate. To verify property 3., note that $X^{\bar{d}}$ is invertible modulo all $p_{l}(X)$. Hence, $\textbf{z}_{ij}=0\mod p_{l}(X)$ implies $\textbf{z}_{ij}^{(h)}=0\mod p_{l}(X)$ for all $l\not\in T$. Moreover,
$\underline{\textbf{z}}_{ij}^{(h)}=X^{-\bar{d}}\overline{\textbf{z}}_{ij}^{(h)}\mod p_{l}(X),l\not\in{\cal T},$
where $X^{-\bar{d}}$ denotes the multiplicative inverse of $X^{\bar{d}}$ modulo $p_{l}(X)$. Since the $p_{l}(X)$ are pairwise co-prime, their product is $\bar{p}(X)$, and $\deg(\bar{p}(X))=\bar{d}>\deg(\underline{\textbf{z}}_{ij}^{(h)})$, property 3. follows from the Chinese Remainder Theorem.
@@
In some special cases, we can do better than Lemma 30, i.e. we can avoid the term $n/g_{\cal T}$. The next lemma considers the case where $n$ is even and $d=2^{e},e\geq 0,$ is a divisor of $n$, which implies that the polynomial $X^{d}+1$ divides $X^{n}+1$
Lemma 31.
Let $r>0$ be such that $r\sqrt{q}<q/2$. Let $n=2^{e}n^{\prime},e\geq 1,n^{\prime}$ odd. Set $d=2^{e}$ and $p_{\cal T}(X)=X^{d}+1$ for some appropriate subset $\cal T$} of the indices of the divisors of $X^{n}+1$ modulo $q$ (see Equation (17)) . Then
$\left|B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}\right|\leq\sigma_{2kd}\left(\frac{d}{n}\right)^{dk}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd}q^{kd}.$ (20)
Proof.
By assumption
$X^{n}+1=(X^{d}+1)(X^{d(n^{\prime}-1)}-X^{d(n^{\prime}-2)}+\cdots-X^{d}+1),$
i.e. in this case
$\bar{p}_{\cal T}(X)=X^{d(n^{\prime}-1)}-X^{d(n^{\prime}-2)}+\cdots-X^{d}+1.$
Let $\textbf{z}=(\textbf{z}_{11},\ldots,\textbf{z}_{1k},\textbf{z}_{21},\ldots,\textbf{z}_{2k})\in B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}$. Then there are elements $\textbf{a}_{ij}\in R,1\leq i\leq 2,1\leq j\leq k$, such that as a polynomial element $\textbf{a}_{ij}$ has degree at most d1d-1 and for all index pairs $(i,j)$ we have
$\textbf{z}_{ij}=\textbf{a}_{ij}\bar{p}_{\cal T}\mod qR.$
Because of the special structure of $\bar{p}_{\cal T}(X)$ we get
$\textbf{a}_{ij}\bar{p}_{\cal T}(X)=\textbf{a}_{ij}X^{d(n^{\prime}-1)}-\textbf{a}_{ij}X^{d(n^{\prime}-2)}+\cdots-\textbf{a}_{ij}X^{d}+\textbf{a}_{ij}\mod qR.$
Since $r\sqrt{q}<q/2$, and as in the proof of Lemma 30, we can ignore the modulo $qR$ and restrict ourselves to elements $\textbf{a}_{ij}$ with coefficients less than $r\sqrt{q}<q/2$. Then from the above we conclude that for $\textbf{z}\in B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}$ we need that
$\|\textbf{a}\|=\left(\sum_{i,j}\|\textbf{a}_{ij}\|^{2}\right)^{1/2}\leq\sqrt{\frac{d}{n}}r\sqrt{q}.$
Since $\textbf{a}\in\mathbb{Z}^{2kd}$ (as a vector), the lemma follows from Lemma 7. ∎
@@
We split the sum into subsets corresponding to subsets $\cal T$$\subset{\cal I}$.
$\sum_{\textbf{s}\in B_{2kn}(r\sqrt{q})\cap\mathbb{Z}^{2kn}\setminus\{0\}}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]=\\ \sum_{Z_{\cal T}\subset{\cal I}}\sum_{\textbf{s}\in Z_{\cal T}\cap B_{2kn}(r\sqrt{q})}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}].$ (22)
@@
Corollary 33.
With the notation from above
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ \sum_{{\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset}\frac{n}{g_{\cal T}}\frac{1}{\sqrt{2\pi d_{\cal T}k}}\left(\frac{\pi e}{nk}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}.$ (23)
@@
For every integer $d$, $\sigma_{2d}=\frac{\pi^{d}}{d!}$. Using Stirling’s formula (Lemma 8), we obtain for every set $\cal T$}
$\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\leq\frac{1}{\sqrt{2\pi d_{\cal T}k}}\left(\frac{\pi e}{d_{\cal T}k}\right)^{kd_{\cal T}}.$
This proves the corollary ∎
Remark 34.
Using Lemma 31, in Theorem 32 as well as in Corollary 33, for all subsets $\cal T$} such that $p_{\cal T}(X)$ is of the form $X^{2^{e}}+1,e\geq 0$, we can drop the factor $n/g_{\cal T}$.

### v2

@@ 8.2 Symplectic lattices from R-SIS lattices
$\bar{I}_{n-1}:=\left(\begin{matrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\vdots&0&0\end{matrix}\right),\qquad\sigma_{n}:=\left(\begin{matrix}1&0\\ 0_{n}&-\bar{I}_{n-1}\end{matrix}\right).$
@@
$M_{\textbf{h}}:=M_{\rho(\textbf{h})}\qquad\text{and}\qquad\bar{M}_{\textbf{h}}:=M_{\bar{\rho}(\textbf{h})}.$
@@ 8.3 Symplectic lattices from M-SIS lattices
that is, we replace entries $\textbf{h}_{ij}$ by the matrix representations of multiplication with $\textbf{h}_{ij}.$
Consider a symmetric matrix H over $R$. Matrix $M_{\rho(\textbf{H})}$ is not symmetric, since matrix representations of multiplications with ring elements are not symmetric. However, for symmetric H the matrix
@@
$M_{\textbf{H}}:=M_{\rho(\textbf{H})}\qquad\text{and}\qquad\bar{M}_{\textbf{H}}:=M_{\bar{\rho}(\textbf{H})}.$
@@ 9.1 Analyzing the probability
$X^{n}+1=\prod_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\prod_{j_{m}\in[n_{m}]}p_{m,j_{m}}(X),$
[…]
${\cal I}:=\bigcup_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\{m\}\times[n_{m}].$
For $\textbf{z}\in R^{2k},\textbf{z}\neq\textbf{0}\mod qR,$ set
$\bar{T}_{m}(\textbf{z}):=\{j\in[n_{m}]:\textbf{z}_{1}=\textbf{z}_{2}=0\mod p_{m,j}(X)\}$ (17)
and
$T_{m}(\textbf{z}):=[n_{m}]\setminus\bar{T}_{m}(\textbf{z}),\qquad T(\textbf{z}):=\bigcup_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\{m\}\times T_{m}(\textbf{z}),\qquad\bar{T}(\textbf{z}):={\cal I}\setminus T(\textbf{z}).$ (18)
Next set
$\ell_{m}({\textbf{z}}):=|T_{m}(\textbf{z})|\qquad\text{and}\qquad p_{\textbf{z}}(X):=\prod_{(m,j_{m})\in T(\textbf{z})}p_{m,j_{m}}(X).$
Consequently,
$d_{\textbf{z}}:=|T(\textbf{z})|=\sum_{\begin{subarray}{c}m\mid 2n\\ m\nmid n\end{subarray}}\ell_{m}(\textbf{z})d_{m}\qquad\text{and}\qquad\deg(p_{\textbf{z}}(X))=d_{\textbf{z}}.$ (19)
@@
$\textbf{z}\in{\cal L}^{\perp}(\textbf{H})\Leftrightarrow\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod qR\Leftrightarrow\forall m,j_{m}\;\;\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{m,j_{m}}(X).$
Since the polynomials $p_{m,j_{m}}(X)$ are irreducible modulo $q$, for each pair $(m,j_{m})$ the equation $\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{m,j_{m}}(X)$ is an equation over the finite field $\mathbb{F}_{q^{d_{m}}}$. For $(m,j_{m})\in\bar{T}(\textbf{z})$ the equation is satisfied for all $\textbf{H}.$ Hence, let $(m,j_{m})\in T(\textbf{z})$. If $\textbf{z}_{1}=\textbf{0}\mod p_{m,j_{m}}(X),$ then $\textbf{z}_{2}=\textbf{0}\mod p_{m,j_{m}}(X),$ which is impossible by definition of $T_{m}(\textbf{z})$. By Lemma 10
$\Pr[\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{m,j_{m}}(X);\textbf{H}\leftarrow\mathbb{F}_{q^{d_{m}}}^{k\times k}\;\text{symmetric}]=\frac{1}{q^{d_{m}}}.$
For distinct pairs $(m,j_{m})$ these events are independent. The lemma follows from the definition of $\ell_{m}(\textbf{z})$ and $d_{\textbf{z}}$. ∎
@@ 9.2 Bounds on the number of short vectors
Let $\cal T$$\subset{\cal I}$. In this subsection we derive bounds on the number of short elements in $\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2})\in R^{2k}$ satisfying $\textbf{z}_{1}=\textbf{z}_{2}=0\mod p_{m,j_{m}}(X)$ exactly for the pairs in $\cal T$}. To state this more precisely, for ${\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset$, set
  • $p_{\cal T}(X):=\prod_{(m,j_{m})\in{\cal T}}p_{m,j_{m}}(X),\qquad\bar{p}_{\cal T}(X):=\prod_{(m,j_{m})\not\in T}p_{m,j_{m}}(X)$
  • $d_{\cal T}:=\deg(p_{\cal T}(X)),\qquad\bar{d}_{\cal T}:=\deg(\bar{p}_{\cal T}(X))$
@@
and
$Z_{\cal T}:=\left\{\textbf{z}\in R^{2k}:\bar{T}(\textbf{z})={\cal T}\right\}.$ (20)
@@
Properties 1. and 2. are immediate. To verify property 3., note that $X^{\bar{d}}$ is invertible modulo all polynomials $p_{m,j_{m}}(X)$. Hence, $\textbf{z}_{ij}=0\mod p_{m,j_{m}}(X)$ implies $\textbf{z}_{ij}^{(h)}=0\mod p_{m,j_{m}}(X)$ for all $(m,j_{m})\not\in T$. Moreover,
$\underline{\textbf{z}}_{ij}^{(h)}=X^{-\bar{d}}\overline{\textbf{z}}_{ij}^{(h)}\mod p_{m,j_{m}}(X),(m,j_{m})\not\in{\cal T},$
where $X^{-\bar{d}}$ denotes the multiplicative inverse of $X^{\bar{d}}$ modulo $p_{m,j_{m}}(X)$. Since the $p_{m,j_{m}}(X)$ are pairwise co-prime, their product is $\bar{p}(X)$, and $\deg(\bar{p}(X))=\bar{d}>\deg(\underline{\textbf{z}}_{ij}^{(h)})$, property 3. follows from the Chinese Remainder Theorem.
@@
9.3 Probability analysis
@@
We split the sum into $Z_{\cal T}$ corresponding to subsets $\cal T$$\subset{\cal I}$.
$\sum_{\textbf{s}\in B_{2kn}(r\sqrt{q})\cap\mathbb{Z}^{2kn}\setminus\{0\}}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]=\\ \sum_{{\cal T}\subset{\cal I}}\sum_{\textbf{s}\in Z_{\cal T}\cap B_{2kn}(r\sqrt{q})}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}].$ (23)
@@
We rewrite the probability estimate above in a form that is better suited to further analysis. To do so, for two sequences $\vec{j}=(j_{1},\ldots,j_{s}),\vec{n}=(n_{1},\ldots,n_{s})$ of natural numbers, we set $\vec{j}\leq\vec{n}$ iff for all $i$ we have $j_{i}\leq n_{i}$. If $\vec{j}$ is not the all zero vector we write $\vec{\neq}\vec{0}$. For $\vec{j}\leq\vec{n}$, define
${\vec{n}\choose\vec{j}}:=\prod_{i=1}^{s}{n_{i}\choose j_{i}}$.
From now on $\vec{n}=(n_{m})_{m\mid 2n,m\nmid n}$ is the sequence of number of factors in the factorization modulo $q$ of the cyclotomic polynomials dividing $X^{n}+1$. Furthermore $(d_{m})_{m\mid 2n,m\nmid n}$ is the sequence of degrees of these factors. For $\vec{j}\leq\vec{n}$ we set
$d_{\vec{j}}:=\sum_{i=1}^{s}j_{i}d_{i}$
and
${\vec{n}\choose\vec{j}}:=\prod_{i=1}^{s}{n_{i}\choose j_{i}}$.
Finally, set
$\alpha:=r+\sqrt{\frac{kn}{2q}}$
Lemma 29.
With the notation above
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ n\sum_{\vec{j}\leq\vec{n},\vec{j}\neq\vec{0}}{\vec{n}\choose\vec{j}}\sigma_{2kd_{\vec{j}}}\left(\frac{d_{\vec{j}}}{n}\right)^{kd_{\vec{j}}}\alpha^{2kd_{\vec{j}}}.$ (24)
@@
For each set $\cal T$} in (22), only $g_{\cal T}$ and $d_{\cal T}$ depend on $\cal T$}. Clearly, $g_{\cal T}$ is lower bounded by $1$. Furthermore, all $\cal T$} that for all $m$ contain the same number $j_{m}$ of elements from $\{m\}\times[n_{m}]$ have the same degree $d_{\vec{j}},\vec{j}=(j_{m})_{m\mid 2n,m\nmid n}$. ∎
Lemma 30.
With the notation from above and with $d:=\min\{d_{m}\}$,
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ \frac{n}{\sqrt{2\pi kd}}\left(\left(1+\left(\frac{e\pi\alpha^{2}}{kn}\right)^{dk}\right)^{\sum n_{m}}-1\right)$ (25)
Proof.
For every integer $d$, $\sigma_{2d}=\frac{\pi^{d}}{d!}$. Using Stirling’s formula (Lemma 8), we obtain for every vector $\vec{j}$
$\sigma_{2kd_{\vec{j}}}\left(\frac{d_{\vec{j}}}{n}\right)^{kd_{\vec{j}}}\leq\frac{1}{\sqrt{2\pi d_{\vec{j}}k}}\left(\frac{\pi e}{d_{\vec{j}}k}\right)^{kd_{\vec{j}}}.$
Next, we can bound the right-hand side in (24) as follows
$n\sum_{\vec{j}\leq\vec{n},\vec{j}\neq\vec{0}}{\vec{n}\choose\vec{j}}\sigma_{2kd_{\vec{j}}}\left(\frac{d_{\vec{j}}}{n}\right)^{kd_{\vec{j}}}\alpha^{2kd_{\vec{j}}}\leq$ $n\sum_{\vec{j}\leq\vec{n},\vec{j}\neq\vec{0}}{\vec{n}\choose\vec{j}}\frac{1}{\sqrt{2\pi d_{\vec{j}}k}}\left(\frac{e\pi}{nk}\right)^{kd_{\vec{j}}}\alpha^{2kd_{\vec{j}}}$
$\leq$ $\frac{n}{\sqrt{2\pi kd}}\sum_{\vec{j}\leq\vec{n},\vec{j}\neq\vec{0}}{\vec{n}\choose\vec{j}}\left(\frac{e\pi}{nk}\right)^{kd_{\vec{j}}}\alpha^{2kd_{\vec{j}}}$
$=$ $\frac{n}{\sqrt{2\pi kd}}\left(\prod_{m}\left(1+\left(\frac{e\pi\alpha^{2}}{kn}\right)^{kd_{m}}\right)^{n_{m}}-1\right)$
$\leq$ $\frac{n}{\sqrt{2\pi kd}}\left(\left(1+\left(\frac{e\pi\alpha^{2}}{kn}\right)^{kd}\right)^{\sum n_{m}}-1\right)$
Theorem 31.
With the notation from above and for every $0<p\leq 1,k\leq n$, if
$r:=\sqrt{\frac{kn}{e\pi}}\left(\frac{p\sqrt{kd}}{n\sum n_{m}}\right)^{\frac{1}{2kd}}-\sqrt{\frac{nk}{2q}}$
then
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq p.$
Proof.
With our choice of $r$, we have
$\alpha:=\sqrt{\frac{kn}{e\pi}}\left(\frac{p\sqrt{kd}}{n\sum n_{m}}\right)^{\frac{1}{2kd}}.$
Then we can bound the right-hand side in (25) as follows, using $(1+\gamma/m)^{m}\leq 1+2\gamma,$ valid for $\gamma\leq 1/2$, and $p\sqrt{kd}/n\leq 1/2$, by choice of $p,k$,
$\frac{n}{\sqrt{2\pi kd}}\left(\left(1+\left(\frac{e\pi\alpha^{2}}{kn}\right)^{kd}\right)^{\sum n_{m}}-1\right)=$ $\frac{n}{\sqrt{2\pi kd}}\left(\left(1+\frac{p\sqrt{kd}}{n\sum n_{m}}\right)^{\sum n_{m}}-1\right)$
$\leq$ $\frac{n}{\sqrt{2\pi kd}}\frac{2p\sqrt{kd}}{n}$
$\leq$ $p.$
9.4 The case $n=2^{e}$
As we are interested not only in asymptotic results but also in exact results with precise constants, and because it is most relevant for the connection of GKP codes to the NTRU cryptosystem, in this section we consider the case that $n$ is a power of two, $n=2^{e}$, and prime $q=5\mod 8$ or $q=3\mod 16$, in more detail.
Let $n=2^{e}$ be a power of $2$. In this case, over the rationals $X^{n}+1=Q_{2m}(X)$ is irreducible (see (11). The proof of Theorem 2’ in Chapter 4 of [20] shows that for $q=5\mod 8,n\geq 8$, the order of $q$ modulo $2n$ is $n/2$. The same proof can be used for $q=3\mod 16$ and $n\geq 16$. By Theorem 2.47 in [22] (see also (12),
$X^{n}+1=p_{2n,1}(X)p_{2n,2}(X),$ (26)
where polynomials $p_{2n,j}(X)$ have degree $n/2$. Using the notation from above we have ${\cal I}=\{(2n,1),(2n,2)\}$, and the possible subsets $\cal T$} in Theorem 28 are
${\cal T}=\{(2n,1),(2n,2)\},\{(2n,1)\},\{(2n,2)\}.$
The corresponding values for $d_{\cal T},g_{\cal T},$ and $n/\gcd(n,g_{\cal T})$ are
$d_{\cal T}=g_{\cal T}=n,n/2,n/2;\qquad n/\gcd(n,g_{\cal T})=1,2,2.$
Taking this into account, as well as $\sum_{m}n_{m}=2,d=\min\{d_{m}\}=n/2$, Lemma 30 can be improved to
Lemma 32.
Let $n=2^{e}$ and $q=5\mod 16$ or $q=3\mod 8$, then
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ \frac{2}{\sqrt{\pi kn}}\left(\left(1+\left(\frac{e\pi\alpha^{2}}{kn}\right)^{kn/2}\right)^{2}-1\right)$ (27)
Using this lemma and some straightforward calculation, we obtain
Theorem 33.
Let $n=2^{e}$ and $q=5\mod 8$ or $q=3\mod 16$. For every $0<p\leq\frac{6}{\sqrt{kn}}$, if
$r:=\sqrt{\frac{kn}{e\pi}}\left(\frac{p\sqrt{kn}}{4}\right)^{\frac{1}{kn}}-\sqrt{\frac{nk}{2q}}$
then
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq p.$